# Rigidity for Patterson–Sullivan systems with applications to random walks and entropy rigidity

### v2

@@ 6. Mixed Shadows and a Shadow Lemma
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$[M_{1}\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z_{1}$
@@
$[M_{2}\smallsetminus\rho(\gamma_{n})^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\rho(\gamma_{n}))]\to Z_{2}$
@@
$[M_{1}\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z_{1}$
@@
$[M_{2}\smallsetminus\rho(\gamma_{n})^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\rho(\gamma_{n}))]\to Z_{2}$
@@ 7. The Main Theorem
$01=\mu_{1}(\Lambda^{\rm con}(\mathscr{H}_{1}))=\mu_{1}\left(\Gamma\cdot\bigcup_{R>0}\bigcap_{n\geq 1}\Lambda_{R}(\mathscr{H}_{1}(n))\right)>$ 0
@@
[…]
@@
$[M_{1}\smallsetminus\tilde{\gamma}_{j,n}^{-1}\operatorname{\mathcal{O}}_{R_{j}}(\tilde{\gamma}_{j,n})]\rightarrow Z_{j}\quad\text{and}\quad[M_{2}\smallsetminus\rho(\tilde{\gamma}_{j,n})^{-1}\operatorname{\mathcal{O}}_{R_{j}}(\rho(\tilde{\gamma}_{j,n}))]\rightarrow Z_{j}^{\prime}$
@@ 8. Convergence groups and expanding coarse-cocycles
To verify Property (PS3) and Property (PS5), assume $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\to+\infty$, and $[M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z$ with respect to the Hausdorff distance. Then $Z$ must be singleton or empty. Then, since $\Gamma$ is a non-elementary convergence group, Property (PS3) and Property (PS5) are true. ∎
@@ 8.2. Measurable isomorphisms
00 0 $=\lim_{n\rightarrow\infty}\frac{1}{\mu(\operatorname{\mathcal{O}}_{R_{j}}(\gamma_{n}))}\mu\left(\left\{y\in\operatorname{\mathcal{O}}_{R_{j}}(\gamma_{n}):\operatorname{d}(f(x),f(y))>\epsilon\right\}\right)$
@@
[…]
@@ 9. Discrete subgroups of Lie groups
[…]
@@ 9.2. Limit sets
For a proof see [LO23, Lem. 2.9] (for $\theta=\Delta$), [KOW23, Lem. 2.4], [CZZ24, Prop. 2.3], or [KLP17, Sect. 4].
@@ 9.3. Transverse subgroups
A $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma<\operatorname{\mathsf{G}}$ is called non-elementary if #$\#\Lambda_{\theta\cup\mathrm{i}(\theta)}(\Gamma)>2$.
@@
It is easy to see that for a $\operatorname{\mathsf{P}}_{\theta}$-transverse $\Gamma<\operatorname{\mathsf{G}}$, the canonical projection $\Lambda_{\theta\cup\mathrm{i}(\theta)}(\Gamma)\to\Lambda_{\theta}(\Gamma)$ is a $\Gamma$-equivariant homeomorphism (cf. [KOW23, Lem. 9.5]). An important feature of a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma<\operatorname{\mathsf{G}}$ is that the $\Gamma$-action on $\Lambda_{\theta}(\Gamma)$ is a convergence action ([KLP17, Thm. 4.16], [CZZ24, Prop. 2.8]) and that there is a natural class of expanding cocycles.
@@
    [KOW23] If $\Gamma$ is Zariski dense and $\sum_{\gamma\in\Gamma}e^{-\delta_{\phi}(\Gamma)\phi(\kappa(\gamma))}=+\infty$, then any $\phi$-PS measure for $\Gamma$ of dimension $\delta_{\phi}(\Gamma)$ is supported on $\Lambda_{\theta}(\Gamma)$.
9.4. Anosov and relatively Anosov groups
@@ 9.5. Irreducible subgroups
Since $B_{\theta}$ is continuous and $\mathcal{F}_{\theta}$ is compact, Property (PS1) holds. Property (PS2) follows from Lemma 9.9. We now show Property (PS3).
Suppose $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, and $[M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance. Since
@@ 9.6. Zariski dense discrete subgroups
By Theorem 9.11, $(\mathcal{F},\Gamma,\sigma_{\phi},\mu)$ is a PS-system. To see Property (PS5), let $\{\gamma_{n}\}\subset\Gamma$ and $R_{n}\rightarrow+\infty$ be sequences so that $[M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance. Since
@@
The set $\Lambda^{\rm con}(\Gamma_{u,r})=\Lambda^{\rm con}(\mathscr{H})$ above is related to the notion of “$u$-directional limit set” discussed in [Lin06, BLLO23, Sam24, KOW25]. When $\Gamma$ is an irreducible lattice and $\mu$ is a $\mathsf{K}$-invariant measure on $\mathcal{F}$, it follows from the work of Link [Lin06] that $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ for all large $r>0$. For general $\Gamma$ and $\mu$, it was shown by Burger–Landesberg–Lee–Oh [BLLO23] that $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ holds for large $r>0$ if and only if the right-multiplication of $\exp(u\mathbb{R})$ on $\Gamma\backslash\operatorname{\mathsf{G}}/\mathsf{M}$ is ergodic with respect to a Bowen–Margulis–Sullivan measure associated to $\mu$ (see also [KOW25]). It was also shown in [BLLO23] that if $\Gamma<\operatorname{\mathsf{G}}$ is $\operatorname{\mathsf{P}}_{\Delta}$-Anosov and $\operatorname{rank}G\leq 3$, $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ for some $u\in\operatorname{int}\mathfrak{a}^{+}$ and all large $r>0$.
@@ 9.7. Tukia’s theorem in higher rank
9.8. The Linear Case
For use in Section 13 we specialize some of the above discussion to the case when $\operatorname{\mathsf{G}}=\operatorname{\mathsf{PGL}}(d,\operatorname{\mathbb{R}})$. In this case, we can let
$\mathfrak{a}=\{{\rm diag}(a_{1},\dots,a_{d}):a_{1}+\cdots+a_{d}=0\}$
and
$\mathfrak{a}^{+}=\{{\rm diag}(a_{1},\dots,a_{d})\in\mathfrak{a}:a_{1}\geq\cdots\geq a_{d}\}.$
Then the Cartan and Jordan projections are given by
$\kappa(g)=(\log\sigma_{1}(g),\dots,\log\sigma_{d}(g))\quad\text{and}\quad\lambda(g)=(\log\lambda_{1}(g),\dots,\log\lambda_{d}(g))$
where $\sigma_{1}(g)\geq\cdots\geq\sigma_{d}(g)$ are the singular values and $\lambda_{1}(g)\geq\cdots\geq\lambda_{d}(g)$ are the absolute values of the generalized eigenvalues of some (any) representative of $g$ in $\operatorname{\mathsf{GL}}(d,\operatorname{\mathbb{R}})$ with determinant ±$\pm 1$.
With this choice of $\mathfrak{a}^{+}$, $\Delta=\{\alpha_{1},\dots,\alpha_{d-1}\}$ where
$\alpha_{j}({\rm diag}(a_{1},\dots,a_{d}))=a_{j}-a_{j+1}$
and the opposition involution satisfies $\mathrm{i}(\alpha_{j})=\alpha_{d-j}$.
We also let $\omega_{j}\in\mathfrak{a}^{*}$ denote the fundamental weight associated to $\alpha_{j}$, which satisfies
$\omega_{j}({\rm diag}(a_{1},\dots,a_{d}))=a_{1}+\cdots+a_{j}.$
Notice that when $\theta\subset\Delta$,
$\mathfrak{a}_{\theta}^{*}=\left\langle\omega_{j}|_{\mathfrak{a}_{\theta}}:\alpha_{j}\in\theta\right\rangle.$
Given $\theta=\{\alpha_{j_{1}},\dots,\alpha_{j_{k}}\}\subset\Delta$ with $j_{1}<\cdots<j_{k}$, the parabolic subgroup $\operatorname{\mathsf{P}}_{\theta}$ is the stabilizer of the partial flag
$\left\langle e_{1},\dots,e_{j_{1}}\right\rangle\subset\left\langle e_{1},\dots,e_{j_{2}}\right\rangle\subset\cdots\subset\left\langle e_{1},\dots,e_{j_{k}}\right\rangle$
where $e_{1},\dots,e_{d}$ is the standard basis of $\operatorname{\mathbb{R}}^{d}$. So we can identify $\operatorname{\mathcal{F}}_{\theta}$ with the partial flag manifold $\operatorname{\mathcal{F}}_{j_{1},\dots,j_{k}}(\operatorname{\mathbb{R}}^{d})$ and $\operatorname{\mathcal{F}}_{\mathrm{i}(\theta)}$ with the partial flag manifold $\operatorname{\mathcal{F}}_{d-j_{k},\dots,d-j_{1}}(\operatorname{\mathbb{R}}^{d})$. Using these identifications, two flags $x=(x^{j_{i}})_{i=1}^{k}\in\operatorname{\mathcal{F}}_{\theta}$ and $y=(y^{d-j_{i}})_{i=1}^{k}\in\operatorname{\mathcal{F}}_{\mathrm{i}(\theta)}$ are transverse if and only if $x^{j_{i}}$ and $y^{d-j_{i}}$ are transverse for all $i=1,\dots,k$.
To avoid cumbersome notation, in this setting we often replace $\theta$ subscripts with the indices appearing in $\theta$, e.g. if $\theta=\{\alpha_{1},\alpha_{d-1}\}$, then
$\operatorname{\mathsf{P}}_{1,d-1}=\operatorname{\mathsf{P}}_{\theta},\quad\operatorname{\mathcal{F}}_{1,d-1}=\operatorname{\mathcal{F}}_{\theta},\quad\text{and}\quad\Lambda_{1,d-1}(\Gamma)=\Lambda_{\theta}(\Gamma).$
The standard inner product on $\operatorname{\mathbb{R}}^{d}$ induces an inner product on $\wedge^{j}\operatorname{\mathbb{R}}^{d}$ where $\{e_{i_{1}}\wedge\cdots\wedge e_{i_{j}}\}$ is an orthonormal basis. Given $v\in\wedge^{j}\operatorname{\mathbb{R}}^{d}$, we let $\left\|v\right\|$ denote the norm induced by this inner product. Then when $\alpha_{j}\in\theta$, the partial Iwasawa cocycle satisfies
(27) $\omega_{j}\left(B_{\theta}(g,x)\right)=\log\frac{\left\|\tilde{g}(v_{1}\wedge\cdots\wedge v_{j})\right\|}{\left\|v_{1}\wedge\cdots\wedge v_{j}\right\|}$
where $x^{j}=\left\langle v_{1},\dots,v_{j}\right\rangle$ and $\tilde{g}$ is some (any) representative of $g$ in $\operatorname{\mathsf{GL}}(d,\operatorname{\mathbb{R}})$ with determinant ±$\pm 1$.
Recall that a subgroup $\Gamma<\operatorname{\mathsf{PGL}}(d,\operatorname{\mathbb{R}})$ is irreducible if there are no $\Gamma$-invariant proper linear subspaces and strongly irreducible if every finite index subgroup is irreducible. We will use the following result of Labourie.
Proposition 9.17.
[Lab06, Prop. 10.3] If $\Gamma<\operatorname{\mathsf{PGL}}(d,\operatorname{\mathbb{R}})$ is strongly irreducible, then $\Gamma$ is strongly $\operatorname{\mathsf{P}}_{\theta}$-irreducible for every non-empty $\theta\subset\Delta$.
@@ 10. Group actions with contracting isometries
Suppose $\Gamma<\operatorname{Mod}(\Sigma)$ is non-elementary and $\mu$ is a Busemann PS-measure for $\Gamma$ of dimension $\delta$ on $\partial_{GM}\operatorname{\mathcal{T}}$. Then $\mu$ is part of a well-behaved PS-system with respect to some hierarchy =$\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$ and with magnitude function $\gamma\mapsto\operatorname{d}_{\operatorname{\mathcal{T}}}(o,\gamma o)$ for a fixed $o\in\operatorname{\mathcal{T}}$. Moreover, if $\sum_{\gamma\in\Gamma}e^{-\delta\operatorname{d}_{\operatorname{\mathcal{T}}}(o,\gamma o)}=+\infty,$ then

### v3

@@ 6. Mixed Shadows and a Shadow Lemma
$h(x)=\lim_{n\to\infty}\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})}h(y)d\mu_{1}(y)$
@@
$0=\lim_{n\to\infty}\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}(\gamma_{n})}|h(y)-h(x)|d\mu_{1}(y).$
@@
$0\leq\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})\right)}$ $\int_{\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})}|h(y)-h(x)|d\mu_{1}(y)$
$\leq\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}(\gamma_{n})}|h(y)-h(x)|d\mu_{1}(y)$
$\leq\frac{C}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}(\gamma_{n})}|h(y)-h(x)|d\mu_{1}(y)\to 0.$
@@
$\left|h(x)-\lim_{n\to\infty}\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})}h(y)d\mu_{1}(y)\right|$
$\qquad\qquad\leq\lim_{n\to\infty}\frac{1}{\mu_{1}\left(\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})\right)}\int_{\operatorname{\mathcal{O}}_{R}^{f}(\gamma_{n})}|h(y)-h(x)|d\mu_{1}(y)=0.$
@@
$M_{1}\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z_{1}$
@@
$M_{2}\smallsetminus\rho(\gamma_{n})^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\rho(\gamma_{n}))]\to Z_{2}$
@@
$M_{1}\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z_{1}$
@@
$M_{2}\smallsetminus\rho(\gamma_{n})^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\rho(\gamma_{n}))]\to Z_{2}$
@@ 7. The Main Theorem
$1=\mu_{1}(\Lambda^{\rm con}(\mathscr{H}_{1}))=\mu_{1}\left(\Gamma\cdot\bigcup_{R>0}\bigcap_{n\geq 1}\Lambda_{R}(\mathscr{H}_{1}(n))\right)>0$
@@
$h(x_{0})=\lim_{n\rightarrow\infty}\frac{1}{\mu_{1}(\operatorname{\mathcal{O}}^{f}_{R}(\gamma_{n}))}\int_{\operatorname{\mathcal{O}}^{f}_{R}(\gamma_{n})}hd\mu_{1}=\lim_{n\rightarrow\infty}\frac{\tilde{\mu}_{2}(\operatorname{\mathcal{O}}^{f}_{R}(\gamma_{n}))}{\mu_{1}(\operatorname{\mathcal{O}}^{f}_{R}(\gamma_{n}))}.$
@@
$M_{1}\smallsetminus\tilde{\gamma}_{j,n}^{-1}\operatorname{\mathcal{O}}_{R_{j}}(\tilde{\gamma}_{j,n})]\rightarrow Z_{j}\quad\text{and}\quad[M_{2}\smallsetminus\rho(\tilde{\gamma}_{j,n})^{-1}\operatorname{\mathcal{O}}_{R_{j}}(\rho(\tilde{\gamma}_{j,n}))]\rightarrow Z_{j}^{\prime}$
@@ 8. Convergence groups and expanding coarse-cocycles
To verify Property (PS3) and Property (PS5), assume $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\to+\infty$, and $M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z$ with respect to the Hausdorff distance. Then $Z$ must be singleton or empty. Then, since $\Gamma$ is a non-elementary convergence group, Property (PS3) and Property (PS5) are true. ∎
@@ 8.2. Measurable isomorphisms
$0$ $=\lim_{n\rightarrow\infty}\frac{1}{\mu(\operatorname{\mathcal{O}}_{R_{j}}(\gamma_{n}))}\mu\left(\left\{y\in\operatorname{\mathcal{O}}_{R_{j}}(\gamma_{n}):\operatorname{d}(f(x),f(y))>\epsilon\right\}\right)$
@@
$\frac{1}{C_{j}}e^{-\delta\left\|\gamma_{n}\right\|}\leq\frac{d{\gamma_{n}^{-1}}_{*}\mu}{d\mu}\leq C_{j}e^{-\delta\left\|\gamma_{n}\right\|}\quad\mu\text{-a.e.}$
@@ 9. Discrete subgroups of Lie groups
Let $X:=\operatorname{\mathsf{G}}/\mathsf{K}$ and fix a basepoint $o=[e]\in\operatorname{\mathsf{G}}/\mathsf{K}$. Fix a $\mathsf{K}$-invariant norm $\|\cdot\|$ on $\mathfrak{a}$ induced from the Killing form, and let $\operatorname{d}_{X}$ denote the $\operatorname{\mathsf{G}}$-invariant symmetric Riemannian metric on $X$ defined by $\operatorname{d}_{X}(go,ho)=\|\kappa(g^{-1}h)\|$ for $g,h\in\operatorname{\mathsf{G}}$.
@@ 9.2. Limit sets
For a proof see [LO23, Lem. 2.9] (for $\theta=\Delta$), [KOW25b, Lem. 2.4], [CZZ24, Prop. 2.3], or [KLP17, Sect. 4].
@@ 9.3. Transverse subgroups
A $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma<\operatorname{\mathsf{G}}$ is called non-elementary if $\Lambda_{\theta\cup\mathrm{i}(\theta)}(\Gamma)>2$.
@@
It is easy to see that for a $\operatorname{\mathsf{P}}_{\theta}$-transverse $\Gamma<\operatorname{\mathsf{G}}$, the canonical projection $\Lambda_{\theta\cup\mathrm{i}(\theta)}(\Gamma)\to\Lambda_{\theta}(\Gamma)$ is a $\Gamma$-equivariant homeomorphism (cf. [KOW25b, Lem. 9.5]). An important feature of a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma<\operatorname{\mathsf{G}}$ is that the $\Gamma$-action on $\Lambda_{\theta}(\Gamma)$ is a convergence action ([KLP17, Thm. 4.16], [CZZ24, Prop. 2.8]) and that there is a natural class of expanding cocycles.
@@
    [KOW25b] If $\Gamma$ is Zariski dense and $\sum_{\gamma\in\Gamma}e^{-\delta_{\phi}(\Gamma)\phi(\kappa(\gamma))}=+\infty$, then any $\phi$-PS measure for $\Gamma$ of dimension $\delta_{\phi}(\Gamma)$ is supported on $\Lambda_{\theta}(\Gamma)$.
9.4. (relatively) Anosov groups
@@ 9.5. Irreducible subgroups
Since $B_{\theta}$ is continuous and $\mathcal{F}_{\theta}$ is compact, Property (PS1) holds. Property (PS2) follows from Lemma $9.$ We now show Property (PS3).
Suppose $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, and $M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance. Since
@@ 9.6. Zariski dense discrete subgroups
By Theorem 9.11, $(\mathcal{F},\Gamma,\sigma_{\phi},\mu)$ is a PS-system. To see Property (PS5), let $\{\gamma_{n}\}\subset\Gamma$ and $R_{n}\rightarrow+\infty$ be sequences so that $M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance. Since
@@
The set $\Lambda^{\rm con}(\Gamma_{u,r})=\Lambda^{\rm con}(\mathscr{H})$ above is related to the notion of “$u$-directional limit set” discussed in [Lin06, BLLO23, Sam24, KOW25a]. When $\Gamma$ is an irreducible lattice and $\mu$ is a $\mathsf{K}$-invariant measure on $\mathcal{F}$, it follows from the work of Link [Lin06] that $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ for all large $r>0$. For general $\Gamma$ and $\mu$, it was shown by Burger–Landesberg–Lee–Oh [BLLO23] that $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ holds for large $r>0$ if and only if the right-multiplication of $\exp(u\mathbb{R})$ on $\Gamma\backslash\operatorname{\mathsf{G}}/\mathsf{M}$ is ergodic with respect to a Bowen–Margulis–Sullivan measure associated to $\mu$ (see also [KOW25a]). It was also shown in [BLLO23] that if $\Gamma<\operatorname{\mathsf{G}}$ is $\operatorname{\mathsf{P}}_{\Delta}$-Anosov and $\operatorname{rank}G\leq 3$, $\mu(\Lambda^{\rm con}(\Gamma_{u,r}))=1$ for some $u\in\operatorname{int}\mathfrak{a}^{+}$ and all large $r>0$.
@@ 10. Group actions with contracting isometries
Suppose $\Gamma<\operatorname{Mod}(\Sigma)$ is non-elementary and $\mu$ is a Busemann PS-measure for $\Gamma$ of dimension $\delta$ on $\partial_{GM}\operatorname{\mathcal{T}}$. Then $\mu$ is part of a well-behaved PS-system with respect to some hierarchy $\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$ and with magnitude function $\gamma\mapsto\operatorname{d}_{\operatorname{\mathcal{T}}}(o,\gamma o)$ for a fixed $o\in\operatorname{\mathcal{T}}$. Moreover, if $\sum_{\gamma\in\Gamma}e^{-\delta\operatorname{d}_{\operatorname{\mathcal{T}}}(o,\gamma o)}=+\infty,$ then